OpenStudy (anonymous):

Find the slope and y -intercept of 6x-5y=-30.

OpenStudy (kropot72):

Rearrange the terms to the standard form of a linear equation: y = mx + c Where m is the slope and c is the y intercept. 6x - 5y = -30 -5y = -30 - 6x 5y = 30 + 6x 5y = 6x + 30 Divide both sides by 5: y = (6/5)x + 6 Therefore the slope is 6/5 = 1.2 and the y intercept is 6

OpenStudy (anonymous):

when you do the equation you get y=6(x+5)/5 you then do distribute property n you get y=6x+30/5 you then divide so you'll be left with y=6/5x+6 and we know y=mx+b is slope intercept so slope would be 6/5 and y intercet= (0.6)